OpenStudy (anonymous):

A juggler throws a ball into the air from a height of 5 ft with an initial vertical velocity of 16 ft/s. The function that models the height h of the ball in feet t seconds after the ball is thrown is h(t) = -16t2 + 16t + 5. How long does the juggler have to catch the ball before it hits the ground? _______ s.

OpenStudy (imstuck):

Solve this by factoring it. When h = 0, the ball is at the ground. So you have to find where the function = 0 in order to find out how long he has before the ball hits the ground. When you solve a quadratic formula by factoring, you are in actuality finding where the function = 0, and that's what we are looking for here: the time it will take for the height of the ball to be 0.

OpenStudy (imstuck):

So let's factor!

OpenStudy (imstuck):

I'm assuming you are familiar with the quadratic formula?

OpenStudy (anonymous):

-(4t-5)(4t+1)=0

OpenStudy (imstuck):

That's right! Which one of those is your solution, then? Remember that time cannot be a negative number.

OpenStudy (anonymous):

zero is -1.5 and 1.5 so answer is 1.5?

OpenStudy (imstuck):

Actually, it's this...follow me for a sec...

OpenStudy (anonymous):

no 1.45 sorry

OpenStudy (anonymous):

made a mistake

OpenStudy (imstuck):

(4t - 5) = 0...4t = 5...t = 5/4...t = 1.25 seconds (not 1.5)

OpenStudy (anonymous):

tq

OpenStudy (imstuck):

is that a "ty"? I think it is! You're welcome! Thank you for the medal!